# Enhancing Multi-Agent Collaboration with Attention-Based Actor-Critic Policies

### v2

@@ Abstract
This paper introduces Team-Attention-Actor-Critic (TAAC), a reinforcement learning algorithm designed to enhance multi-agent collaboration in cooperative environments. TAAC employs a Centralized Training/Centralized Execution scheme incorporating multi-headed attention mechanisms in both the actor and critic. This design facilitates dynamic, inter-agent communication, allowing agents to explicitly query teammates, thereby efficiently managing the exponential growth of joint-action spaces while ensuring a high degree of collaboration. We further introduce a penalized loss function which promotes diverse yet complementary roles among agents. We evaluate TAAC in a simulated soccer environment against benchmark algorithms representing other multi-agent paradigms, including Proximal Policy Optimization and Multi-Agent Actor-Attention-Critic. We find that TAAC exhibits superior performance and enhanced collaborative behaviors across a variety of metrics (win rates, goal differentials, Elo ratings, inter-agent connectivity, balanced spatial distributions, and frequent tactical interactions such as ball possession swaps).
Index Terms:
Reinforcement Learning, Multi-Agent Systems, Actor-Critic, Attention Mechanism, Centralized Training and Execution (CTE)
@@ I Introduction
The best methods for training a team to collaborate in a cooperative environment remains an open question. The main paradigms that have been used to learn collaboration across agents in cooperative MARL are Decentralized Training/Decentralized Execution (DTDE), Centralized Training/Decentralized Execution (CTDE), and Centralized Training/Centralized Execution (CTCE). In DTDE agents operate independently without shared information both during training and inference while in CTDE information is shared during training (usually to estimate values) but not when agents select actions or during inference. Finally, CTCE shares information during execution, which allows delegation of control to a single controller which determines the joint-action of all agents. DTDE and CTDE schemes present challenges due to the need to decide whether and how to share information across agents; these modeling choices are critical for team collaboration. [^1] Decentralized execution schemes can also suffer from convergence problems due to the inherent non-stationarity that arises when multiple agents learn and adapt concurrently. The CTCE single controller avoids these difficulties but poses another problem: the full joint-action space under a CTCE scheme is the product set of individual agent action sets and so its size grows exponentially in the number of agents. Effective team collaboration would seem to hinge on a robust search of the joint-action space for complimentary actions across agents, and so a CTCE scheme imposes scalability constraints on collaborative activity.
To address these challenges, we introduce and test an RL algorithm that leverages shared information to foster inter-agent collaboration. Our proposed model, Team-Attention-Actor-Critic (TAAC), is a CTCE scheme that incorporates an attention mechanism directly into the actor to model an information-sharing “conversation” between multiple agents which enables them to dynamically query teammates during decision-making. This allows our agents to manage the joint-action space in a scalable fashion. We further introduce a loss that encourages agents to learn diverse yet complementary roles, enhancing collaboration. We evaluate TAAC in a simulated soccer environment, an ideal testbed for teamwork and spatial coordination, and benchmark its performance against PPO (as an example of DTDE) and MAAC (as an example of CTDE) using both standard metrics (e.g., win rate, Elo) and multiple indicators of collaboration (e.g., pairwise distance, possession swaps, connectivity).
@@ II Related Work
In cooperative MARL environments, CTCE seems a natural choice to maximize collaborative potential, especially given the robustness of modern information sharing capabilities which would seem to allow for relatively costless message sending where agents either do not wish to or cannot restrict information sharing. A number of recent papers on MARL have proposed managing agent collaboration by leveraging modern deep learning architectures. Within the CTCE paradigm, [1] and [2] break down the full joint-action space by requiring agents to select actions sequentially and treating MARL as a sequence modeling problem by applying transformers. [3] and [4] apply attention to deal with high-dimensional action spaces in a value-based context (to variants of deep $Q$-learning). By contrast, we focus on simultaneous action selection in a policy-based algorithmic context. 222[1] and [2] would converge to a Markov Perfect Equilibrium for a stochastic game with within-period sequential actions by the agents; there are no guarantees that the equilibrium set for such a game would map well onto the equilibrium set for a game with true simultaneous choices.
@@
Beyond this, some CTDE algorithms have applied attention to model communication between agents, particularly [5] and [6] which use attention mechanisms to learn when and with whom to communicate. The former models communication groups dynamically via attention gating. The latter models targeted communication using attention mechanisms, enabling message exchange between specific agents. Finally and perhaps most relevant for our work, MAAC [7] learns a centralized critic with a soft attention mechanism that dynamically selects agents to attend to at each time step during training.  [^3] Other CTDE schemes have explored different approaches to deal with information within multi-agent environments by modeling communication and information sharing between agents via a combination of heuristic rules (e.g. [8], [9]), and explicitly modeling message selection (e.g. [10]), message timing (e.g. [11] extending [12]), or message recipients (e.g. [13], [14]).
@@
These CTDE methods are particularly appropriate for teaching agents to manage environments collaboratively in which information sending is problematic, e.g. because agents are physically separated, because messages are costly to send, or because there are incentives to maintain private information. At the same time, if information sending is not the main constraint  [^4] Examples of such environments might be sports teams, simulated environments, digital games, or any environment in which a slight time lag to build a full picture of observations across agents is acceptable. then the focus on decentralized execution leads to more complicated methods than might be needed  [^5] Predefined or hueristic communication architectures restrict communication and hence restrain potential collaboration among agents. A potential critique of [8] and [9] is that use of an aggregative operation across agents has the potential to eliminate differences among neighbouring agents and result in loss of important information. , likely underestimates the extent to which collaborative action is possible, and relegates it to no more than an implicit, emergent property during inference.
@@
In this paper, we build upon [7] and [15]. We extend their work by incorporating an attention mechanism in the actor and a novel loss which allows agents to receive information in a selective manner. This allows us to let agents collaborate explicitly during inference. It functions by allowing an agent to pose an open-ended question to its peers. For example, in a soccer game scenario, a player with the ball may query, ‘Who should I pass to?’ Agents which are not nearby opponents may be identified as optimal recipients, and their responses inform the passing decision. We additionally develop a novel loss (named conformity loss) which encourages diverse actions in agents, enhancing collaboration in certain environments. We have named this model Team-Attention-Actor-Critic (TAAC).
To test our algorithm, we train it in a simulation of a soccer environment. In this setting, agents control individual players on a team competing in a virtual soccer match. We present our results by comparing the performance of our approach against state-of-the-art algorithms using a combination of traditional performance metrics (such as win rates and ELO ratings) and collaboration metrics (such as inter-agent connectivity and pairwise distance).
This paper is organized as follows. In Section 3, we introduce the notation and provide the necessary background. Section 4 details our proposed algorithm along with its implementation. In Section 5, we describe the experimental setup and the evaluation metrics employed. Section 6 presents the experimental results, while Section 7 offers an in-depth analysis and discussion of these findings. Finally, Section 8 concludes the paper and outlines directions for future work.
@@ III Background and preliminaries
Our strategy to model collaboration will build upon actor-critic algorithms. Below, we introduce this algorithm along with the principles and notation needed to formalize it. Following this section, we provide a complete description of our architecture.
In reinforcement learning, control problems with multiple agents can be modelled as Markov Decision Games (MDG), which are defined by:
@@
[…]
$\vec{z}_{\backslash i}=\langle z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n}\rangle$
and we will write $\vec{z}_{i}$ to mean the pair $(z_{i},\vec{z}_{\backslash i})$.
In environments that take place over a finite number of discrete periods $T$, the sequence of periods in which the agents participate is referred to as an episode. At each time step, each agent makes an observation, $o_{i,t}=O_{i}(s_{t})$, of the global state, which encapsulates all available information for agent $i$; the vector of such observations across agents is $\vec{o}_{t}=\langle o_{i,t}\rangle_{i=1,\ldots,n}$ The goal of each agent is to learn a policy which describes the probability that a trained agent should take action $a_{i}$ given an observation $o_{i,t}$, to maximize its individual sequence of rewards throughout an episode: $\sum_{t=0}^{T-1}\gamma^{t}r_{i}(s_{t+1},\vec{a}_{t},s_{t})$. Here $\vec{a}_{t}$ and $s_{t}$ are the agents’ actions and the common state at time $t$ and $\gamma\in[0,1]$ is the discount factor on future rewards. In general, policies could differ across agents; we will focus on the case where the policy is shared. Any differences in decision-making across agents are modelled using their observations of the state. Given this, we write these policies as $\pi(a_{i}|o_{i,t})$.
@@ III-A Policy Gradients
Policy gradient methods model this policy directly, usually via a function approximator with parameters $\mathbf{u}$; we indicate the dependence of the policy on these parameters by writing $\pi_{\mathbf{u}}(\cdot)$. [^6] By contrast, value-based methods (e.g. $Q$-learning, [16]) estimate the state-conditional value of actions and then use those values to build a policy. To learn $\mathbf{u}$ for $\pi_{\mathbf{u}}(\cdot)$, these methods use an estimate of the policy gradient ([17], [18], [19]) to update the parameters in the direction of increase in the expected sum of discounted rewards. In the case of $n$ agents with shared policies, taking the Monte Carlo approach of using the full return, the policy gradient may be written as:
@@
Finally, $b(\cdot)$ is a baseline subtracted from the return to avoid high variance in the estimates of $\nabla_{\mathbf{u}}J(\mathbf{u})$ across episodes driven by use of $G_{i,t}(\tau)$. [^7] An arbitrary function of the state can be introduced into the policy gradient equation without changing the expectation – see [17] for details. If $b(\cdot)$ is allowed to depend on agent actions then this will bias the estimate of the gradient.
@@ III-B Actor-Critic Algorithms
A common choice for the baseline function $b(\cdot)$ is an estimate of the long-term value, for example, the state-value function, $V(o_{i,t})$, which predicts the return to an agent given an observation $o_{i,t}$ (see [20]). If we model the state value function as a differentiable function of parameters $\mathbf{w}$ and write $V_{\mathbf{w}}(\cdot)$ then this yields the following baseline:
$b(o_{i,t})$ $=V_{\mathbf{w}}(o_{i,t})$
$=\sum_{a\in A_{i}}\pi_{\mathbf{u}}(a|o_{i,t})Q_{\mathbf{w}}(o_{i,t},a)$
where the second equality follows from the usual definitions of the state-value function $V_{\mathbf{w}}(\cdot)$ and the state-action value function $Q_{\mathbf{w}}(\cdot)$. We will use this form of the baseline below to model multi-agent structure.
The policy is referred to as the actor and the long-term value estimate as the critic, and the resulting model is called an actor-critic algorithm. The actor’s parameters are learned through the policy gradient estimates and the critic’s parameters are learned through value-based methods (e.g. gradient descent on the difference between the critic’s predicted state value and a target value, which can be either the observed $G_{i,t}(\tau)$ or a bootstrapped value, see [21]).
@@ IV-A Attention Mechanism
We extend [7] by introducing a separate attention mechanism in the actor. The intuition behind doing this is to allow agents to use the internal representations of other agents directly when making choices, with the idea being that enabling this is analogous to allowing agents to “put themselves into the shoes” of their colleagues and thereby giving them insight into what other agents will choose. If they build an understanding of what their colleagues will choose then they may be able to create collaboration by selecting actions that dovetail well.
In general, attention works as follows (see [22]). An initial input $z$ is suitably embedded as a sequence of vectors $M(z)=\{m_{1}(z),\ldots,m_{n}(z)\}$ and multiplied with learned matrices $W_{Q}$, $W_{K}$, and $W_{V}$. Attention is then computed as:
@@
where $W=(W_{Q},W_{K},W_{V})$ and $d_{K}$ is the dimensionality of the $M(z)W_{K}$ matrix. [^8] The subscripts refer to the usual presentation of $W_{Q}$, $W_{K}$, and $W_{V}$ as the query, key, and value matrices, respectively. Note that in all other sections of the paper $Q(\cdot)$ will refer to the state-action value function. Division by $d_{K}$ is done to scale the values for numerical stability.
In a natural language application, the input might be a sequence of embedded tokens (e.g. words). In our application, these embedded input tokens will represent individual agent observation and/or action information. We apply multiple attention heads in parallel, each with its own set of weight matrices. The updated outputted sequence of tokens from these heads are concatenated by agent to create an updated set of agent-specific embedding information. Specifically, given $h$ attention heads, for agent $i$ we gather updated agent-specific embedding information by concatenating:
@@ IV-B Actor Attention Architecture
    Each updated agent-specific embedding is passed through a set MLP layers, say $m^{\pi}_{2}(\cdot|u_{m^{\pi}_{2}})$ where $u_{m_{2}^{\pi}}$ are associated parameters, and a final softmax operation is applied to generate $n$ distributions over actions, one for each agent.
@@ IV-C Critic Attention Architecture
In the critic, we incorporate both the observations and actions of other agents. Consequently, the critic’s state-action values are written as $Q_{\mathbf{w}}(\vec{o},\vec{a})\in\mathbb{R}^{n}$ so that $Q_{\mathbf{w}}(\vec{o},\vec{a})=\langle Q^{(1)}_{\mathbf{w}}(\vec{o},\vec{a}),\ldots,Q^{(n)}_{\mathbf{w}}(\vec{o},\vec{a})\rangle$ and $Q^{(i)}_{\mathbf{w}}(\vec{o},\vec{a})$ provides that state-action value estimate for agent $i$ taking action $a_{i}\in A_{i}$. The forward pass of the parameterized critic network is as follows:
  1. 1.
    For $i=1,\ldots,n$ the observation $o_{i}$ and action $a_{i}$ are passed through an appropriate embedding layer, say $m_{1}^{Q}(\cdot|w_{m_{1}^{Q}})$ where $w_{m_{1}^{Q}}$ are associated parameters, to generate initial embeddings which are formed into the matrix:
    $M_{Q}(\vec{o},\vec{a}|w_{m_{1}^{Q}})=\left(\begin{array}[]{c}m^{c}_{1}(o_{1},a_{1}|w_{m_{1}^{Q}})\\ \vdots\\ m^{c}_{1}(o_{n},a_{n}|w_{m_{1}^{Q}})\end{array}\right).$
  2. 2.
    These embeddings are passed into a layer of $h_{Q}$ attention heads which yields:101010in [7], their attention mechanism differs only by having the $V$ matrix be fixed for each attention head.
    $A_{j}(M_{c}(\vec{o},\vec{a}|w_{m_{1}^{Q}})|w_{A_{j}})$
    where $w_{A_{j}}=(w_{A_{j},K},w_{A_{j},Q},w_{A_{j},V})$ are the parameters for the $j$th critic attention head.
  3. 3.
    For every agent $i=1,\ldots,n$ we gather updated agent-specific embedding information by concatenating as described in equation 4, which we write in terms of the input observation and action as $E_{i}(\vec{o},\vec{a}|w_{m_{1}^{Q}},w_{A})$ where $w_{A}=(w_{A_{1}},\ldots,w_{A_{h_{Q}}})$.
  4. 4.
    Each updated agent-specific embedding is combined with the original embedding and passed into a set of MLP layers, say $m^{Q}_{2}(\cdot|w_{m_{2}^{Q}})$ where $w_{m_{2}^{Q}}$ are associated parameters, to generate $n$ state-action value estimates, one for each agent.
With all these steps in mind we may write the output of the policy for the $i$th agent as:
$\pi^{(i)}_{\mathbf{u}}(\cdot|\vec{o})=\mbox{Softmax}\left(m^{a}_{2}\left(\left.E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\right|u_{m_{2}^{\pi}}\right)\right)$
where $\mathbf{u}=(u_{m_{1}^{\pi}},u_{A},u_{m_{2}^{\pi}})$ and the output of the critic for the $i$th agent:
@@
where $\mathbf{w}=(w_{m_{1}^{q}},w_{A},w_{m_{2}^{q}})$, respectively.
We conclude the description of the architecture with two notes. First, we deliberately omitted the original embedding as a direct input to the post-attention layers of the actor. This design choice forces the actor to rely solely on the information provided by the attention blocks rather than bypassing them, ensuring collaborative policies. By contrast, in the critic architecture, we include the original embedding to provide additional information, as the critic does not require the same enforced reliance on collaborative signals as the actor does. Second, each of the $n$ outputs in both the actor and critic make use of all model parameters, meaning that there is no obvious way to factor either model into agent-specific policies and critics that do not share parameters and information. The actor and critics architectures are depicted in figure  1 and  2, respectively.

### v3

@@ Abstract
This paper introduces Team-Attention-Actor-Critic (TAAC), a reinforcement learning algorithm designed to enhance multi-agent collaboration in cooperative environments. TAAC employs a Centralized Training/Centralized Execution scheme incorporating multi-headed attention mechanisms in both the actor and critic. This design facilitates dynamic, inter-agent communication, allowing agents to explicitly query teammates, thereby efficiently managing the exponential growth of joint-action spaces while ensuring a high degree of collaboration. We evaluate TAAC in a diverse range of cooperative environments and benchmark its performance against other multi-agent paradigms, including Proximal Policy Optimization (PPO) and Multi-Actor-Attention-Critic (MAAC). We find that TAAC outperforms other algorithms as the level of coordination and collaboration required increases and that it performs comparably to the other algorithms in environments which require less coordination and collaboration.
publicationid: pubid: This is an extended version of the paper published in the Proceedings of the 2025 IEEE International Conference on Machine Learning and Applications (ICMLA 2025).
@@ I Introduction
The best methods for training a team to collaborate in a cooperative environment remain an open question. The main paradigms that have been used to learn collaboration across agents in cooperative MARL are Decentralized Training/Decentralized Execution (DTDE), Centralized Training/Decentralized Execution (CTDE), and Centralized Training/Centralized Execution (CTCE). In DTDE, agents operate independently without shared information both during training and inference while in CTDE, information is shared during training (usually to estimate values) but not when agents select actions or during inference. Finally, CTCE shares information during execution, which allows delegation of control to a single controller that determines the joint-action of all agents. DTDE and CTDE schemes present challenges due to the need to decide whether and how to share information across agents; these modeling choices are critical for team collaboration. [^1] Decentralized execution schemes can also suffer from convergence problems due to the inherent non-stationarity that arises when multiple agents learn and adapt concurrently. The CTCE single controller avoids these difficulties but poses another problem: the full joint-action space under a CTCE scheme is the product set of individual agent action sets and so its size grows exponentially in the number of agents. Effective team collaboration would seem to hinge on a robust search of the joint-action space for complementary actions across agents, and so a CTCE scheme imposes scalability constraints on collaborative activity.
To address these challenges, we introduce and test a MARL algorithm that leverages shared information to foster inter-agent collaboration. Our proposed model, Team-Attention-Actor-Critic (TAAC), is a CTCE scheme that incorporates an attention mechanism directly into the actor to model an information-sharing “conversation" between multiple agents which enables them to dynamically query teammates during decision-making. This allows our agents to manage the joint-action space in a scalable fashion. We evaluate TAAC in a diverse set of collaborative environments and benchmark its performance against PPO (as an example of DTDE) and MAAC (as an example of CTDE).
@@ II Related Work
In cooperative MARL environments, CTCE seems a natural choice to maximize collaborative potential, especially given the robustness of modern information sharing capabilities which would seem to allow for relatively costless message sending where agents either do not wish to or cannot restrict information sharing. A number of recent papers on MARL have proposed managing agent collaboration by leveraging modern deep learning architectures. Within the CTCE paradigm, [Wen2022] and [Liu2024] break down the full joint-action space by requiring agents to select actions sequentially and treating MARL as a sequence modeling problem by applying transformers. [Tavakoli2019] and [info15050279] apply attention to deal with high-dimensional action spaces in a value-based context (to variants of deep $Q$-learning). By contrast, we focus on simultaneous action selection in a policy-based algorithmic context.222[Wen2022] and [Liu2024] would converge to a Markov Perfect Equilibrium for a stochastic game with within-period sequential actions by the agents; there are no guarantees that the equilibrium set for such a game would map well onto the equilibrium set for a game with true simultaneous choices.
@@
Numerous CTDE schemes have explored different approaches to deal with information within multi-agent environments by modeling communication and information sharing between agents via a combination of heuristic rules (e.g. [yang2020meanfieldmultiagentreinforcement], [sukhbaatar2016learningmultiagentcommunicationbackpropagation]), and explicitly modeling message selection (e.g. [foerster2016learningcommunicatedeepmultiagent]), message timing (e.g. [DOC] extending [optioncritic]), or message recipients (e.g. [liu2019multiagentgameabstractiongraph], [jiang2020graphconvolutionalreinforcementlearning]). Beyond these, some CTDE algorithms have applied attention to model communication between agents, particularly [jiang2018learningattentionalcommunicationmultiagent] and [das2020tarmactargetedmultiagentcommunication] which use attention mechanisms to learn when and with whom to communicate. The former models communication groups dynamically via attention gating. The latter models targeted communication using attention mechanisms, enabling message exchange between specific agents. Finally and perhaps most relevant for our work, MAAC [MAAC] learns a centralized critic with a soft attention mechanism that dynamically selects agents to attend to at each time step during training.
@@
These CTDE methods are particularly appropriate for teaching agents to manage environments collaboratively in which information sending is problematic, e.g. because agents are physically separated, because messages are costly to send, or because there are incentives to maintain private information. At the same time, if information sending is not the main constraint [^3] Examples of such environments might be sports teams, simulated environments, digital games, or any environment in which a slight time lag to build a full picture of observations across agents is acceptable. then the focus on decentralized execution leads to more complicated methods than might be needed [^4] Predefined or heuristic communication architectures restrict communication and hence restrain potential collaboration among agents. A potential critique of [yang2020meanfieldmultiagentreinforcement] and [sukhbaatar2016learningmultiagentcommunicationbackpropagation] is that use of an aggregative operation across agents has the potential to eliminate differences among neighboring agents and result in loss of important information., likely underestimates the extent to which collaborative action is possible, and relegates it to no more than an implicit, emergent property during inference.
@@
In this paper, we build upon [MAAC] and [COMA]. We extend their work by incorporating an attention mechanism in the actor which allows agents to receive information in a selective manner. This allows us to let agents collaborate explicitly during inference. It functions by allowing an agent to pose an open-ended question to its peers. For example, in a soccer game scenario, a player with the ball may query, ‘Who should I pass to?’ Agents which are close to the goal may be identified as optimal recipients, and their responses inform the passing decision and ultimately improve collaboration. We have named this model Team-Attention-Actor-Critic (TAAC).
@@ III Background and preliminaries
In reinforcement learning, control problems with multiple agents can be modeled as a Partially observable Markov game [PartiallyObservableStochasticGames], which is defined by:
@@
In environments that take place over a finite number of discrete periods $T$, the sequence of periods in which the agents participate is referred to as an episode. At each time step, each agent makes an observation, $o_{i,t}=O_{i}(s_{t})$, of the global state, which encapsulates all available information for agent $i$; the vector of such observations across agents is $\vec{o}_{t}=\langle o_{i,t}\rangle_{i=1,\ldots,n}$. The goal of each agent is to learn a policy which describes the probability that a trained agent should take action $a_{i}$ given an observation $o_{i,t}$, to maximize its individual sequence of rewards throughout an episode: $\sum_{t=0}^{T-1}\gamma^{t}r_{i}(s_{t+1},\vec{a}_{t},s_{t})$. Here $\vec{a}_{t}$ and $s_{t}$ are the agents’ actions and the common state at time $t$ and $\gamma\in[0,1]$ is the discount factor on future rewards. In general, policies could differ across agents; we will focus on the case where the policy is shared. Any differences in decision-making across agents are modeled using their observations of the state. Given this, we write these policies as $\pi(a_{i}|o_{i,t})$.
@@ III-A Policy Gradients
Policy gradient methods model this policy directly, usually via a function approximator with parameters $\mathbf{u}$; we indicate the dependence of the policy on these parameters by writing $\pi_{\mathbf{u}}(\cdot)$. To learn $\mathbf{u}$ for $\pi_{\mathbf{u}}(\cdot)$, these methods use an estimate of the policy gradient ([PolicyGradientTheorem], [williams-1988], [williams-1992]) to update the parameters in the direction of increase in the expected sum of discounted rewards. In the case of $n$ agents with shared policies, taking the Monte Carlo approach of using the full return, the policy gradient may be written as:
@@
Finally, $b(\cdot)$ is a baseline subtracted from the return to avoid high variance in the estimates of $\nabla_{\mathbf{u}}J(\mathbf{u})$ across episodes driven by use of $G_{i,t}(\tau)$. [^5] An arbitrary function of the state can be introduced into the policy gradient equation without changing the expectation – see [PolicyGradientTheorem] for details. If $b(\cdot)$ is allowed to depend on agent actions then this will bias the estimate of the gradient.
@@ III-B Actor-Critic Algorithms
A common choice for the baseline function $b(\cdot)$ is an estimate of the long-term value, for example, the state-value function, $V(o_{i,t})$, which predicts the return to an agent given an observation $o_{i,t}$ (see [suttonbarto2018]). If we model the state value function as a differentiable function of parameters $\mathbf{w}$ and write $V_{\mathbf{w}}(\cdot)$ then this yields the following baseline:
$b(o_{i,t})$ $=V_{\mathbf{w}}(o_{i,t})=\sum_{a\in A_{i}}\pi_{\mathbf{u}}(a|o_{i,t})Q_{\mathbf{w}}(o_{i,t},a)$
where the second equality follows from the usual definitions of the state-value function $V_{\mathbf{w}}(\cdot)$ and the state-action value function $Q_{\mathbf{w}}(\cdot)$. [^6] We use the Counterfactual Multi-agent baseline described in [COMA] in the actors’ policy gradients.
The policy is referred to as the actor and the long-term value estimate as the critic, and the resulting model is called an actor-critic algorithm. The actor’s parameters are learned through the policy gradient estimates and the critic’s parameters are learned through value-based methods (e.g. gradient descent on the difference between the critic’s predicted state value and a target value, which can be either the observed $G_{i,t}(\tau)$ or a bootstrapped value, see [mnih2013]).
@@ IV-A Attention Mechanism
We extend [MAAC] by introducing a separate attention mechanism in the actor. The intuition behind doing this is to allow agents to use the internal representations of other agents directly when making choices, with the idea that enabling this is analogous to allowing agents to “put themselves into the shoes" of their colleagues and thereby giving them insight into what other agents will choose. If they build an understanding of what their colleagues will choose, then they may be able to create collaboration by selecting actions that dovetail well.
In general, attention works as follows (see [attentionneed]). An initial input $z$ is suitably embedded as a sequence of vectors $M(z)=\{m_{1}(z),\ldots,m_{n}(z)\}$ and multiplied with learned matrices $W_{Q}$, $W_{K}$, and $W_{V}$. Attention is then computed as:
@@
where $W=(W_{Q},W_{K},W_{V})$ and $d_{K}$ is the dimensionality of the $M(z)W_{K}$ matrix. [^7] The subscripts refer to the usual presentation of $W_{Q}$, $W_{K}$, and $W_{V}$ as the query, key, and value matrices, respectively. Note that in all other sections of the paper $Q(\cdot)$ will refer to the state-action value function. [^8] Division by $d_{K}$ is done to scale the values for numerical stability.
In a natural language application, the input might be a sequence of embedded tokens (e.g. words). In our application, these embedded input tokens will represent individual agent observation and/or action information. We apply multiple attention heads in parallel, each with its own set of weight matrices. The updated output sequence of tokens from these heads are concatenated by agent to create an updated set of agent-specific embedding information. Specifically, given $h$ attention heads, for agent $i$ we gather updated agent-specific embedding information by concatenating:
@@ IV-B Actor Attention Architecture
    Each updated agent-specific embedding is passed through a set of MLP layers, say $m^{\pi}_{2}(\cdot|u_{m^{\pi}_{2}})$ where $u_{m_{2}^{\pi}}$ are associated parameters, and a final softmax operation is applied to generate $n$ distributions over actions, one for each agent.
With all these steps in mind, we may write the output of the policy for the $i$th agent as:
$\pi^{(i)}_{\mathbf{u}}(\cdot|\vec{o})=\mbox{Softmax}\left(m^{a}_{2}\left(m^{a}_{1}(o|u_{m_{1}^{\pi}}),\left.E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\right|u_{m_{2}^{\pi}}\right)\right)$
where $\mathbf{u}=(u_{m_{1}^{\pi}},u_{A},u_{m_{2}^{\pi}})$.
@@ IV-C Critic Attention Architecture
In the critic, we incorporate both the observations and actions of other agents. Consequently, the critic’s state-action values are written as $Q_{\mathbf{w}}(\vec{o},\vec{a})\in\mathbb{R}^{n}$ so that $Q_{\mathbf{w}}(\vec{o},\vec{a})=\langle Q^{(1)}_{\mathbf{w}}(\vec{o},\vec{a}),\ldots,Q^{(n)}_{\mathbf{w}}(\vec{o},\vec{a})\rangle$ and $Q^{(i)}_{\mathbf{w}}(\vec{o},\vec{a})$ provides that state-action value estimate for agent $i$ taking action $a_{i}\in A_{i}$. The forward pass of the parameterized critic network is analogous to that used for the policy; the initial embedding depends on parameters $w_{m_{1}^{Q}}$, there are $h_{Q}$ attention heads with $w_{A_{j}}=(w_{A_{j},K},w_{A_{j},Q},w_{A_{j},V})$ the parameters for the $j$th critic attention head, the post-attention output is concatenated by agent as $E_{i}(\vec{o},\vec{a}|w_{m_{1}^{Q}},w_{A})$ where $w_{A}=(w_{A_{1}},\ldots,w_{A_{h_{Q}}})$, and the output MLP layer uses parameters $w_{m_{2}^{Q}}$ to generate $n$ state-action value estimates, one for each agent. The output of the critic for the $i$th agent is:
@@
where $\mathbf{w}=(w_{m_{1}^{Q}},w_{A},w_{m_{2}^{Q}})$.
The actor and critic architectures are depicted in Figure 1 and Figure 2, respectively.